OpenStudy (anonymous):

Help (one medal) Find the side length of a square with the given area. 121 in²

ganeshie8 (ganeshie8):

\(\large A = s^2 \)

ganeshie8 (ganeshie8):

so, \(\large s = \sqrt{A} = \sqrt{121} = ?\)

OpenStudy (texaschic101):

you figure it out yet ?

OpenStudy (anonymous):

Good job gane